# HiSTF Mamba: Hierarchical Spatiotemporal Fusion with Multi-Granular Body-Spatial Modeling for High-Fidelity Text-to-Motion Generation

### v1

@@ 3.4 Dynamic Spatiotemporal Fusion Module
$\mathbf{T}$ $=\frac{\text{Softmax}(QK^{T})}{\sqrt{N/r}}\,V,$
$Z$ $=W_{2}\cdot\text{Concat}\bigl{(}S,\;T\bigr{)},$
@@
3.5 Loss
@@
4 Experiments
4.1 Datasets and Evaluation Metrics
@@
Five evaluation metrics from [20] are adopted, including (1) Frechet Inception Distance (FID), which serves as the primary metric to assess the difference in feature distributions between generated motions and real motions; (2) R-Precision evaluates how well a generated motion aligns with its corresponding textual description. It measures the retrieval accuracy by checking whether the ground truth description ranks among the top-k closest matches based on Euclidean distances between motion and text features; (3) MultiModal Distance (MM Dist), which calculates the average distance between generated motions and their corresponding text descriptions; (4) Diversity measures the variety of generated motions and is computed as the variance of motion features; (5) Multimodality evaluates the variety of motions generated from the same text description.
4.2 Comparative Studies
We evaluated our method on the HumanML3D [20] and KIT-ML [46] datasets and compared it with several existing state-of-the-art approaches. Specifically, we adopted the network structure described in Section 3.2, setting the number of model layers to $N=3$ and $L=2$. Meanwhile, to accelerate the sampling process, we employed DMP-solver++ and selected $step=10$ and $step=15$ for inference. As shown in Tables 1 and 2, our method achieves significant performance improvements on both the HumanML3D and KIT-ML datasets. On HumanML3D, when $step=10$ is used, our FID metric is further reduced from 0.281 (obtained by the state-of-the-art Motion Mamba) to 0.189, representing an improvement of more than 30%, and we also observe increased R-Precision and MultiModal Distance. On KIT-ML, our method similarly outperforms existing approaches in terms of R-Precision, FID, and MM Dist, with the MM Distance reduced to 2.819.
4.3 Ablation Studies
We investigate how the number of Bi-Temporal Mamba blocks ($N$), the total number of layers ($L$), and the presence of key components: DSFM and part-based learning in Dual-Spatial Mamba) affect model performance. Table 3 provides the results on the HumanML3D dataset with an inference step of 10. The complete HiSTF Mamba configuration, which includes parallel Part-based and Whole-based learning in Dual-Spatial Mamba, Bi-Temporal Mamba, the DSFM module, and uses $N=3$ with $L=2$, achieves the best FID of 0.189. By contrast, merely increasing $N$ to 4 or $L$ to 3 does not consistently improve performance; the FID ranges from 0.303 to 0.397. This finding indicates that simply increasing the number of Bi-Temporal Mamba modules or stacking more spatio-temporal layers, the FID metric does not steadily decrease; instead, it increases or fluctuates within a certain range. This observation indicates that merely stacking spatiotemporal modules does not effectively improve generation quality. In contrast, removing the DSFM module leads to a significant increase in FID form 0.189 to 0.319, showing that the effective fusion of spatiotemporal features provided by DSFM is crucial for achieving higher-quality motion generation. e part-based parallel strategy from Dual-Spatial Mamba and omitting the DSFM module. Discarding part-based learning raises the FID to 0.349, while removing DSFM increases it to 0.319.Furthermore, eliminating part-based modeling in Dual-Spatial Mamba also increases the FID from 0.189 to 0.349, suggesting that losing the ability to learn fine-grained joint details reduces the quality of the generated motions.
In summary, the performance improvement of HiSTF Mamba primarily arises from the Dual-Spatial Mamba’s local-global parallel spatial modeling and the DSFM module’s dynamic fusion of spatiotemporal features. These two components work in synergy to maintain overall motion coherence, capture the fine-grained details of human joints, effectively fuse spatiotemporal features, and ultimately achieve a significantly lower FID.
Table 3: Ablation study on HumanML3D with inference step=10. We vary $N$, $L$, remove part-based learning in Dual-Spatial Mamba, and remove the DSFM module. For simplicity, we use BTM to represent Bi-Temporal Mamba, and P-b and W-b to denote the part-based learning and whole-based learning in Dual-Spatial Mamba, respectively.
BTM W-b P-b DSFM N L FID
2 2 0.313
3 2 0.189
4 2 0.303
2 3 0.335
3 3 0.397
4 3 0.363
3 2 0.349
3 2 0.319
4.4 Implementation Details
All models in the HiSTF Mamba use the AdamW optimizer, with the learning rate maintained at $1\times 10^{-4}$. The average training step is about 570k, and we set the batch size to 64. We fix the diffusion step count at 1000 and use 10 or 15 sampling steps. On a single NVIDIA GEFORCE RTX 4090 GPU, training on the HumanML3D dataset takes approximately 37.5 hours. The average inference time is about 0.69 seconds.
5 Conclusion
In this paper, we have introduced HiSTF Mamba, a novel method for text-to-motion generation that integrates local and global spatial features with bidirectional temporal modeling. Our Dual-Spatial Mamba block effectively captures fine-grained joint dynamics alongside coarse whole-body movements, while the Bi-Temporal Mamba block encodes both short-term and long-term dependencies in a bidirectional manner. We also introduced the DSFM block to perform dynamic fusion of refined temporal features with spatial representations, thereby reducing redundancy and enhancing overall spatiotemporal expressiveness.
Extensive experiments on HumanML3D and KIT-ML demonstrate that HiSTF Mamba not only reduces FID by nearly 30% in HumanML3D relative to state-of-the-art methods, but also delivers high-fidelity and semantically aligned motion sequences. In addition, our method achieves these gains with efficient training times and rapid inference. We believe HiSTF Mamba offers a promising approach for generating more realistic, detail-rich human motions, and serves as a foundation for future work in text-to-motion generation.

### v2

@@ 3.4 Dynamic Spatiotemporal Fusion Module
$\mathbf{T}$ $=\frac{\text{Softmax}(QK^{T})}{\sqrt{N/r}}\,V,\quad Z=W_{2}\cdot\text{Concat}\bigl{(}S,T\bigr{)}.$
@@ 4 Experiments
Table 2: Comparison of quantitative results on the KIT-ML dataset
4.1 Datasets
Figure 3: Qualitative comparison of typical methods. We randomly selected three textual samples to generate motion sequences. Our results exhibit closer alignment with the text and present richer joint-level details.
4.2 Comparative Studies
We evaluated our method on the HumanML3D [28] and KIT-ML [61] datasets and compared it with several existing state-of-the-art approaches. Our method models motion data in a continuous representation, whereas existing VQ-VAE-based approaches discretize motion into tokens. Since each token typically corresponds to a short motion segment already present in the dataset, this discretization provides a natural advantage when computing the FID metric: the generated motions are more likely to resemble the real data distribution, resulting in better FID scores. In light of this, we do not conduct a direct comparison with current VQ-VAE-based state-of-the-art methods [36, 37, 38, 57, 62]. Specifically, we adopted the network structure described in Section 3.1, setting the number of model layers to $N=3$ and $L=2$. Meanwhile, to accelerate the sampling process, we employed DPM-Solver++ and selected $step=10$ and $step=15$ for inference. As shown in Tables 1 and 2, our method achieves significant performance improvements on both the HumanML3D and KIT-ML datasets. On HumanML3D, when $step=10$ is used, our FID metric is further reduced from 0.281 (obtained by the state-of-the-art Motion Mamba) to 0.189, representing an improvement of more than 30%, and we also observe increased R-Precision and MultiModal Distance. On KIT-ML, our method similarly outperforms existing approaches in terms of R-Precision, FID, and MM Dist, with the MM Distance reduced to 2.819. Additionally, we present qualitative motion results in Fig. 3. Compared with other approaches, our method produces motions that more closely align with the text, exhibiting more detailed joint movements and better full-body coordination.
4.3 Ablation Studies
In this experiment, we investigate how the number of Bi-Temporal Mamba blocks ($N$), the total number of layers ($L$), and the presence or absence of key components DSFM, (whole-based learning, and part-based learning within Dual-Spatial Mamba’s two branches), as well as the entire Dual-Spatial Mamba and Bi-Temporal Mamba, affect the model’s performance on the HumanML3D dataset with an inference step of 10.
Ablation of layer. As shown in Table 3, we investigate how varying the number of Bi-Temporal Mamba blocks ($N$) and the total number of layers ($L$) impacts model performance. When configured with $N=3$ and $L=2$, our approach achieves the lowest Fréchet Inception Distance (FID) of $0.189$, indicating a better capture of the underlying motion distribution. Although increasing $N$ to 3 leads to gains in Multimodality, it also increases the parameter count and inference time, making $L=2$ and $N=3$ a favorable choice for balancing accuracy and computational cost.
Ablation of key components. Table 4 presents the ablation results for each key component. When ablating the entire Dual-Spatial Mamba and Bi-Temporal Mamba, we replace them with the same number of standard Transformer blocks for a fair comparison. Although this substitution slightly increases the number of parameters, it raises the FID to 0.367 and 0.439, indicating that removing these modules forfeits the ability to capture both global body coordination and fine-grained limb details. Moreover, ablating DSFM leads to a 207% increase in FID while saving 1.38M parameters, demonstrating that DSFM is effective spatiotemporal fusion with acceptable parameter cost. We also examine the roles of Whole-based and Part-based branches in Dual-Spatial Mamba. Removing the Part-based branch lowers the AITS to 0.18s, showing that the additional latency primarily arises from large matrix operations rather than increased parameter count. Nevertheless, the Part-based branch indeed improves performance, underscoring its effectiveness. These findings confirm the effectiveness of all proposed components in capturing both fine-grained joint details and overall body coordination, thereby enhancing the quality and accuracy of text-to-motion generation.
N L R Precision↑ FID↓ MM Dist↓ Diversity→ MModality↑ Parameters AITS↓
Top 1 Top 2 Top 3
Real 0.511±.003 0.703±.003 0.797±.002 0.002±.000 2.974±.008 9.503±.065 - - -
2 2 $0.484^{\pm.007}$ $0.679^{\pm.006}$ $0.780^{\pm.006}$ $0.358^{\pm.033}$ $3.132^{\pm.016}$ $\textbf{9.693}^{\pm.083}$ $2.530^{\pm.063}$ 11.90M 0.26s
3 2 $\textbf{0.488}^{\pm.005}$ $\textbf{0.685}^{\pm.004}$ $\textbf{0.784}^{\pm.005}$ $\textbf{0.189}^{\pm.018}$ $\textbf{3.101}^{\pm.022}$ $9.712^{\pm.090}$ $2.529^{\pm.044}$ 15.53M 0.28s
4 2 $0.487^{\pm.006}$ $0.675^{\pm.005}$ $0.773^{\pm.005}$ $0.409^{\pm.036}$ $3.223^{\pm.031}$ $9.938^{\pm.066}$ $\textbf{2.622}^{\pm.035}$ 19.17M 0.33s
2 3 $0.469^{\pm.005}$ $0.666^{\pm.006}$ $0.767^{\pm.006}$ $0.490^{\pm.036}$ $3.242^{\pm.024}$ $10.060^{\pm.010}$ $2.619^{\pm.078}$ 17.32M 0.37s
3 3 $0.471^{\pm.005}$ $0.665^{\pm.005}$ $0.767^{\pm.005}$ $0.339^{\pm.036}$ $3.241^{\pm.023}$ $9.783^{\pm.095}$ $2.721^{\pm.056}$ 22.77M 0.39s
4 3 $0.471^{\pm.007}$ $0.667^{\pm.008}$ $0.763^{\pm.007}$ $0.337^{\pm.040}$ $3.298^{\pm.031}$ $9.831^{\pm.085}$ $2.780^{\pm.078}$ 28.22M 0.49s
Table 3: Ablation study on varying the number of Bi-Temporal Mamba blocks ($N$) and the total number of layers, evaluated on HumanML3D with an inference step of 10. AITS is the abbreviation of average inference time per sentence. Parameters refers to the total number of model parameters.
BTM Whole-b Part-b DSFM R Precision↑ FID↓ MM Dist↓ Diversity→ MModality↑ Parameters AITS↓
Top 1 Top 2 Top 3
Real 0.511±.003 0.703±.003 0.797±.002 0.002±.000 2.974±.008 9.503±.065 - - -
$0.482^{\pm.007}$ $0.678^{\pm.006}$ $0.772^{\pm.005}$ $0.368^{\pm.033}$ $3.747^{\pm.029}$ $9.933^{\pm.059}$ $\textbf{2.601}^{\pm.078}$ 14.99M 0.25s
$0.482^{\pm.006}$ $0.677^{\pm.006}$ $0.779^{\pm.005}$ $0.303^{\pm.026}$ $3.177^{\pm.019}$ $9.988^{\pm.092}$ $2.440^{\pm.067}$ 14.43M 0.18s
$0.473^{\pm.006}$ $0.664^{\pm.007}$ $0.764^{\pm.004}$ $0.367^{\pm.035}$ $3.269^{\pm.244}$ $9.847^{\pm.102}$ $2.510^{\pm.104}$ 19.65M 0.21s
$0.478^{\pm.006}$ $0.672^{\pm.007}$ $0.771^{\pm.005}$ $0.439^{\pm.035}$ $3.206^{\pm.019}$ $9.817^{\pm.094}$ $2.536^{\pm.079}$ 23.55M 0.23s
$0.464^{\pm.006}$ $0.661^{\pm.007}$ $0.764^{\pm.004}$ $0.581^{\pm.054}$ $3.198^{\pm.022}$ $\textbf{9.359}^{\pm.099}$ $2.590^{\pm.079}$ 14.15M 0.25s
$\textbf{0.488}^{\pm.005}$ $\textbf{0.685}^{\pm.004}$ $\textbf{0.784}^{\pm.005}$ $\textbf{0.189}^{\pm.018}$ $\textbf{3.101}^{\pm.022}$ $9.712^{\pm.090}$ $2.529^{\pm.044}$ 15.53M 0.28s
Table 4: Ablation study on key components, evaluated on HumanML3D with an inference step of 10. For simplicity, we use BTM to represent Bi-Temporal Mamba, and Part-b and Whole-b to denote the part-based learning and whole-based learning in Dual-Spatial Mamba, respectively.
4.4 Implementation Details
All models in the HiSTF Mamba use the AdamW optimizer, with the learning rate maintained at $1\times 10^{-4}$. The average training step is about 570k, and we set the batch size to 64. We fix the diffusion step count at 1000 and use 10 or 15 sampling steps. On a single NVIDIA GEFORCE RTX 4090 GPU, training on the HumanML3D dataset takes approximately 37.5 hours. The average inference time is about 0.28 seconds, and our model has a total parameter count of 15.53M. Further details on both inference time and total parameters can be found in the appendix A.1.
5 Conclusion
We introduced HiSTF Mamba, a text-to-motion generation framework that combines local and global spatial features with bidirectional temporal modeling. Its Dual-Spatial Mamba block captures both fine-grained joint dynamics and coarse whole-body movements, while the Bi-Temporal Mamba block encodes short-term and long-term dependencies in a bidirectional manner. To fuse refined temporal features with spatial representations, we proposed the DSFM block, which reduces redundancy and boosts overall spatiotemporal expressiveness.
Although our model still has many limitations (appendix A.5), extensive experiments on HumanML3D and KIT-ML confirm that HiSTF Mamba achieves robust performance across multiple evaluation metrics, while producing high-fidelity and semantically aligned motion sequences. Our method achieves these results with efficient training and rapid inference. We believe HiSTF Mamba provides a promising avenue for generating realistic, detail-rich human motions, and it can serve as a foundation for future text-to-motion research.
Appendix A Technical Appendices and Supplementary Material
A.1 Inference Time and Total Parameters
Figure 4: Comparison of HiSTF Mamba’s AITS and FID with other baseline methods.
To compute the Average Inference Time per Sentence (AITS), we follow the procedure in [19] and report AITS in seconds. We measure AITS on the HumanML3D test set [28] with a batch size of 1, excluding any overhead for model and dataset loading. As shown in Fig. 4, our method requires 0.280 s per sentence, which is slower than TEMOS [29] and Motion Mamba [24]. However, this additional runtime yields marked performance gains: our approach reduces the FID to 0.189, outperforming all other methods in the figure. Although our model introduces more complexity at first glance, it maintains a parameter count of just 15.53 M, around 13% fewer than the baseline MDM [60] (17.88 M). In short, our model’s acceptable increase in inference time and complexity yields a improvement in generation quality.
A.2 Loss of HiSTF Mamba
Let $x_{0}$ be the ground-truth motion sequence, $\hat{x}_{0}$ the predicted motion, and $c$ the textual condition. We denote $FK(\cdot)$ as the forward kinematics function, which converts joint rotations into 3D coordinates (or the identity if positions are already provided). The term $N$ indicates the total number of frames in the motion, and $f_{i}$ is the binary mask for foot contact at frame $i$, where $f_{i}\in\{0,1\}$ and is applied only to foot joints in the sequence.
@@ A.2 Loss of HiSTF Mamba
A.3 Evaluation Metrics
@@
In this work, we employ the method described in [28] to evaluate our approach, using five evaluation metrics. including (1) Frechet Inception Distance (FID), which serves as the primary metric to assess the difference in feature distributions between generated motions and real motions; (2) R-Precision evaluates how well a generated motion aligns with its corresponding textual description. It measures the retrieval accuracy by checking whether the ground truth description ranks among the top-k closest matches based on Euclidean distances between motion and text features; (3) MultiModal Distance (MM Dist), which calculates the average distance between generated motions and their corresponding text descriptions; (4) Diversity measures the variety of generated motions and is computed as the variance of motion features; (5) Multimodality evaluates the variety of motions generated from the same text description.
A.4 Details of Whole2Part and Part2Whole
As illustrated in Fig 5, both HumanML3D [28] and KIT [61] skeletons are each divided into six body segments: Root, R_Leg, L_Leg, Backbone, R_Arm, and L_Arm. Each circle represents a specific joint index. Under the part-based scan, joints belonging to the same body segment are grouped together, enabling fine-grained modeling of local kinematic relationships. In contrast, the whole-based scan processes the entire skeleton as a single sequence of joint indices. By integrating both strategies, our Dual-Spatial Mamba simultaneously captures detailed joint-level nuances and preserves global body coordination.
Figure 5: Illustration of the part-based and whole-based skeletal scans on the HumanML3D [28] (left) and KIT [61] (right) datasets.
A.5 Limitations
(1) Matrix operations in part-based learning: Our Dual-Spatial Mamba adopts a Whole2Part and Part2Whole scheme, which involves extensive matrix operations. Specifically, we split the full-body matrix into six distinct body-part matrices and then merge them back. Because the model must continuously learn both fine-grained joint dynamics and overall body coordination, these operations cannot be preprocessed and must be incorporated into the training process, thereby increasing the inference time.
(2) Performance compared to VQ-VAE-based methods: Although our method achieves noticeable performance gains, some evaluation metrics still fall short when compared to state-of-the-art approaches [36, 37, 38, 57, 62] using VQ-VAE.